A scientist wants to make a solution of tribasic sodium phosphate, Na3PO4, for a laboratory experiment. How many grams of Na3PO4 will be needed to produce 550 mL of a solution that has a concentration of Na+ ions of 0.600 M?

mass Na3PO4
Answer in grams
Answer #1
Step1 [Na+]= .600 M ; [ Na3PO4] = (1/3) [Na+] = .200 M

Step2 Moles of Na3PO4 in 550 ml = .200x550/1000 = .11

Step3 Mass of Na3PO4 = .11 x164 = 18.04 g